OpenStudy (anonymous):

(4.5x10^-2)(2.1x10^3)

OpenStudy (anonymous):

\[(\frac{ 45 }{ 10^{3} })(\frac{ 21 }{ 10 } \times 10^3) \] \[\frac{ 45 \times 21 }{ 10 }\] you can solve furthr